OpenStudy (anonymous):

integral dx/sqrt(x^2-9)

OpenStudy (solomonzelman):

Trig substitution

OpenStudy (anonymous):

x=3sec

OpenStudy (solomonzelman):

yes, good!

OpenStudy (solomonzelman):

And dx=?

OpenStudy (anonymous):

dx=3sec(theta)tan(theta)

OpenStudy (solomonzelman):

forgot: d(theta)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \int\frac{1}{\sqrt{(3\sec \theta)^2~-~9}}(3\sec\theta \tan\theta~d\theta) }\)

OpenStudy (solomonzelman):

Stuck?

OpenStudy (anonymous):

3 integra (1/sqrt3(sec(u)^2-9))(tan(u)/cos(u))

OpenStudy (solomonzelman):

You probably know that \(\color{#000000}{ (a\cdot b)^n=a^n\cdot b^n }\), and not just \(\color{#000000}{ \displaystyle a\cdot b^n }\).

OpenStudy (solomonzelman):

it is a hint to where your mistake is.

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \int\frac{1}{\sqrt{(3\sec \theta)^2~-~9}}(3\sec\theta \tan\theta~d\theta) }\) \(\color{#000000}{ \displaystyle 3 \int\frac{\sec\theta \tan\theta}{\sqrt{3^2\sec^2\theta~-~9~}~~}d\theta }\)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle 3 \int\frac{\sec\theta \tan\theta}{\sqrt{9(\sec^2\theta~-1)~}~~}d\theta }\) \(\color{#000000}{ \displaystyle 3 \int\frac{\sec\theta \tan\theta}{3\sqrt{\sec^2\theta-1}~~}d\theta }\) \(\color{#000000}{ \displaystyle \int\frac{\sec\theta \tan\theta}{\sqrt{\sec^2\theta-1}~~}d\theta }\)

OpenStudy (solomonzelman):

continue ....

OpenStudy (solomonzelman):

Yes, you will have to integrate in terms of theta, but then to solve and simplify everything in terms of x...

OpenStudy (solomonzelman):

First finish integrating with \(\theta\).

OpenStudy (anonymous):

got it, thanks for the help

OpenStudy (solomonzelman):

Oh, alright. YW

OpenStudy (anonymous):

i got ln((1/cos(theta))+tan(theta))+C

OpenStudy (priyar):

good! @garmex now write in terms of x..

OpenStudy (anonymous):

hello,\[\ln \left| ((\ x+\sqrt{x^2 -9})/3)\right|\]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

that's what i ended up with

OpenStudy (anonymous):

and thanks for the response

OpenStudy (priyar):

you got it! that's correct! well Done @garmex